# Global Framework for Simultaneous Emulation Across the Nuclear Landscape

### v1

@@ Abstract
We introduce a hierarchical framework that combines ab initio many-body calculations with a Bayesian neural network, developing emulators capable of accurately predicting nuclear properties across the nuclear chart, including multiple isotopes simultaneously. We benchmark our developments using the oxygen isotopic chain, achieving accurate results for ground-state energies and nuclear charge radii, while providing robust uncertainty quantification. Our framework enables global sensitivity analysis of nuclear binding energies and charge radii with respect to the low-energy constants that describe the nuclear force.
@@ Introduction.
Ab initio methods [18, 19, 20, 21, 22, 23, 24, 25] rely on systematically improvable approximations to solve the many-body Schrodinger equation, starting from $\chi$EFT interactions. Despite major advances in recent years, ab initio nuclear calculations of medium and heavy nuclei remain prohibitively expensive due to the exponential scaling of many-body wave functions. While approximate polynomially-scaling methods (e.g. [20, 19]) have been developed, this complexity still imposes stringent resource demands and methodological challenges [26]. Statistical procedures, such as global sensitivity studies, are necessary to evaluate nuclear observables for samples in the range of millions, even for the simplest realistic nuclear interactions  [27]. As a result, surveying the entire nuclear landscape with realistic uncertainty estimation and establishing a connection to the LECs remains an outstanding challenge.
Nuclear emulators have emerged as powerful tools to address these problems, offering efficient approximations of many-body calculations at substantially reduced computational cost and allowing for uncertainty quantification [28]. Techniques such as multi-output multi-fidelity Deep Gaussian Processes [17, 29] and eigenvector continuation [30, 27, 31, 32] have been proposed as surrogate models that approximate the outputs of nuclear calculations with significantly reduced computational costs. However, these methods are so far limited to individual isotopes, restricting their ability to capture correlations and trends across isotopic chains and more globally. Moreover, they still demand costly ab initio calculation train them.
In this work, we introduce BAyesian Neural Network for Atomic Nuclei Emulation (BANNANE), a hierarchical Bayesian neural network (BNN) framework that integrates multi-fidelity datasets with categorical and positional embeddings to emulate nuclear properties globally. This approach introduces a flexible architecture tailored for the prediction of different nuclear properties with vastly reduced computational time. We demonstrate the overarching capabilities of BANNANE by calculating the nuclear binding energies and charge radii of the oxygen isotopic chain, achieving accurate predictions of these observables. Notably, BANNANE enables a global analysis of the sensitivity of these observables to the LECs used in the interactions. Furthermore, we showcase its strong performance in both interpolation and extrapolation, significantly enhancing the predictive power of nuclear emulators.
@@ Results.
We observe that BANNANE correctly reproduces both total binding energies $E_{B}$ and charge radii $R_{ch}$ across the chain with 0.8 MeV and 0.01 fm Root Mean Squared Error (RMSE), respectively. This represents a significant improvement compared to the Eigenvector Continuation-based subspace-projected coupled-cluster [27] which obtains RMSEs of 3 MeV and 0.02 fm for the ground state energies and charge radii respectively in 16O using 128 training points. While BANNANE uses more training points in total for this isotope (715 in total), only 51 samples are of the highest fidelity, making BANNANE much cheaper to train and yet achieving an RMSE of 0.337 MeV and 0.02 fm respectively for the same isotope. Further study of the residuals and benchmark is provided in the supplemental material [38].
Moreover, we note that the discontinuity caused by the shell closure at $N=8$ is still captured by BANNANE. To further study this phenomenon, we performed a dimensional reduction to perform an analysis of the latent space of the model after the attention mechanism using a projection t-distributed Stochastic Neighbor Embedding (t-SNE) dimensionality reduction [39] for visualization. As illustrated in Fig. 3, there is an evident clustering of the isotopes in the $sd$ shell, after the shell closure at $N=8$, whereas the isotopes in the $p$ shell appear to be mapped to distinct clusters, hinting of the model’s ability to capture an underlying nuclear structure.
@@
Figure 5: Emulator-driven convergence of binding energies $E_{B}$ for oxygen isotopes as a function of $e_{\text{max}}$. Solid markers represent experimental results from [40].
@@
Unified Sensitivity Estimation: Beyond providing fast and accurate emulation, BANNANE enables a global sensitivity analysis (GSA) that would be prohibitive with brute-force ab initio methods. With this method, we can probe how much the variance of each LEC (and their covariance) impacts the variance of the observables. In Fig. 6, we apply a variance-based Sobol approach [42] to assess how each LEC affects $E_{B}$ and $R_{ch}$ across oxygen isotopes from 12O to 24O. We plot the main-effect index $S_{i}$ (filled bars), capturing the direct contribution of each LEC to the variance of an observable, and the total-effect index $S_{T}$ (white outlines), accounting for higher-order interactions.
@@ Conclusions.
We have introduced a hierarchical multi-fidelity Bayesian Neural Network for Atomic Nuclei Emulation (BANNANE), which synthesizes low- and high-fidelity datasets across different nuclei. By leveraging learnable embeddings for nucleon numbers and an additive fidelity-specific architecture, BANNANE not only interpolates faithfully among known nuclei, achieving better accuracy than current emulators but also extrapolates with remarkable accuracy to hitherto uncalculated (or unobserved) nuclei. Due to this, BANNANE offers a computationally feasible way to calculate nuclear properties with realistic uncertainties and with an accuracy comparable to those obtained by considerably costly ab initio methods. These developments have enabled us to investigate the overall dependence of nuclear observables on the LECs that govern inter-nucleon interactions. We found that the nuclear binding energy and charge radii of oxygen isotopes exhibit distinct sensitivities to these LECs, with particular interest in data from neutron-deficient oxygen isotopes that have yet to be measured. We hope our findings will motivate charge radius measurements of these isotopes, which are now feasible with the development of the new Facility for Rare Isotope Beams (FRIB). More broadly, we anticipate that BANNANE will serve as a powerful tool to guide future experiments and establish direct connections between measurements of binding energy, charge radii, and specific components of the nuclear force.
@@ Acknowledgements.
We thank A. Ekstrom, C. Forssén, G. Hagen, and W. G. Jiang for providing the interaction samples used in this work, and J. Holt for insightful discussions. The IMSRG code used in this work makes use of the Armadillo C++ library [44, 45]. This work was supported by the Office of Nuclear Physics, U.S. Department of Energy, under grants DESC0021176 and DE-SC0021179. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [PDF-587464-2024].

### v2

# Global Framework for Emulation of Nuclear Calculations

@@ Abstract
We introduce a hierarchical framework that combines ab initio many-body calculations with a Bayesian neural network, developing emulators capable of accurately predicting nuclear properties across isotopic chains simultaneously and being applicable to different regions of the nuclear chart. We benchmark our developments using the oxygen isotopic chain, achieving accurate results for ground-state energies and nuclear charge radii, while providing robust uncertainty quantification. Our framework enables global sensitivity analysis of nuclear binding energies and charge radii with respect to the low-energy constants that describe the nuclear force.
@@ Introduction.
Ab initio methods [18, 19, 20, 21, 22, 23, 24, 25] rely on systematically improvable approximations to solve the many-body Schrödinger equation, starting from $\chi$EFT interactions. Despite major advances in recent years, ab initio nuclear calculations of medium and heavy nuclei remain prohibitively expensive due to the exponential scaling of many-body wave functions. While approximate polynomially-scaling methods (e.g. [20, 19]) have been developed, this complexity still imposes stringent resource demands and methodological challenges [26, 27, 28]. Statistical procedures, such as global sensitivity studies, are necessary to evaluate nuclear observables for samples in the range of millions, even for the simplest realistic nuclear interactions  [29]. As a result, surveying the entire nuclear landscape with realistic uncertainty estimation and establishing a connection to the LECs remains an outstanding challenge.
Nuclear emulators have emerged as powerful tools to address these problems, offering efficient approximations of many-body calculations at substantially reduced computational cost and allowing for uncertainty quantification [30]. Techniques such as multi-output multi-fidelity Deep Gaussian Processes [17, 31], eigenvector continuation [32, 29, 33, 34, 35] and many-more (e.g. [36, 37, 38]) have been proposed as surrogate models that approximate the outputs of nuclear calculations with significantly reduced computational costs. However, these methods are so far limited to individual isotopes, restricting their ability to capture correlations and trends across isotopic chains and more globally. Moreover, they still demand costly ab initio calculation train them.
To overcome the challenges mentioned above, we introduce BAyesian Neural Network for Atomic Nuclei Emulation (BANNANE), a hierarchical Bayesian neural network (BNN) framework that integrates multi-fidelity datasets with categorical and positional embeddings to emulate nuclear properties globally. This approach introduces a flexible architecture tailored for the prediction of different nuclear properties with vastly reduced computational time. We demonstrate the overarching capabilities of BANNANE by calculating the nuclear binding energies and charge radii of the oxygen isotopic chain, achieving accurate predictions of these observables. Notably, BANNANE enables a global analysis of the sensitivity of these observables to the LECs used in the interactions. This allows us to directly study how macroscopic nuclear properties are impacted by specific details of the nuclear force, and how these sensitivities vary along isotopic chains, which, to our knowledge, is not possible with current nuclear emulators.
@@ Results.
We observe that BANNANE correctly reproduces both total binding energies $E_{B}$ and charge radii $R_{ch}$ across the chain with 0.8 MeV and 0.01 fm Root Mean Squared Error (RMSE), respectively. This represents a significant improvement compared to the Eigenvector Continuation-based subspace-projected coupled-cluster [29] which obtains RMSEs of 3 MeV and 0.02 fm for the ground state energies and charge radii respectively in 16O using 128 training points. While BANNANE uses more training points in total for this isotope (715 in total), only 51 samples are of the highest fidelity, making BANNANE much more computationally efficient to train, while still achieving an RMSE of 0.337 MeV and 0.02 fm, respectively, for the same isotope. Further study of the residuals and benchmark is provided in the supplemental material [44].
Moreover, we note that the discontinuity caused by the shell closure at $N=8$ is well captured by BANNANE. To further study this phenomenon, we performed a dimensional reduction to perform an analysis of the latent space of the model after the attention mechanism using a projection t-distributed Stochastic Neighbor Embedding (t-SNE) dimensionality reduction [45] for visualization. As illustrated in Fig. 3, there is an evident clustering of the isotopes in the $sd$ shell, after the shell closure at $N=8$, whereas the isotopes in the $p$ shell appear to be mapped to distinct clusters, hinting of the model’s ability to capture an underlying nuclear structure.
@@
Figure 5: Emulator-driven convergence of binding energies $E_{B}$ for oxygen isotopes as a function of $e_{\text{max}}$. Results are compared to fully converged VS-IMSRG results [46] at $e_{\text{max}}=14$ using the EM(1.8/2.0) nuclear interaction in purple squares. Solid lines represent experimental results from [47].
@@
Unified Sensitivity Estimation: Beyond providing fast and accurate emulation, BANNANE enables global sensitivity analysis (GSA) that would be computationally prohibitive using brute-force ab initio methods. With this method, we can probe how much the variance of each LEC (and their covariance) impacts the variance of the observables. In Fig. 6, we apply a variance-based Sobol approach [49] to assess how each LEC affects $E_{B}$ and $R_{ch}$ across oxygen isotopes from 12O to 24O. We plot the main-effect index $S_{i}$ (filled bars), capturing the direct contribution of each LEC to the variance of an observable, and the total-effect index $S_{T}$ (white outlines), accounting for higher-order interactions (see supplemental material [44] for more details.).
@@ Conclusions.
We have introduced a hierarchical multi-fidelity Bayesian Neural Network for Atomic Nuclei Emulation (BANNANE), which synthesizes low- and high-fidelity datasets across different nuclei. By leveraging learnable embeddings for nucleon numbers and an additive fidelity-specific architecture, BANNANE not only interpolates faithfully among known nuclei, achieving better accuracy than current emulators but also extrapolates with remarkable accuracy to hitherto uncalculated (or unobserved) nuclei. Due to this, BANNANE offers a computationally feasible way to calculate nuclear properties with realistic uncertainties and with an accuracy comparable to those obtained by considerably costly ab initio methods. These developments have enabled us to investigate the overall dependence of nuclear observables on the LECs that govern inter-nucleon interactions. We found that the nuclear binding energy and charge radii of oxygen isotopes exhibit distinct sensitivities to these LECs, with particular interest in data from neutron-deficient oxygen isotopes that have yet to be measured. We hope our findings will motivate charge radius measurements of these isotopes, which are now feasible via laser spectroscopy experiments [50] at the new Facility for Rare Isotope Beams (FRIB) [51]. More broadly, we anticipate that BANNANE will serve as a powerful tool to guide future experiments and establish direct connections between measurements of binding energy, charge radii, and specific components of the nuclear force.
@@ Acknowledgements.
We thank A. Ekström, C. Forssén, G. Hagen, and W. G. Jiang for providing the interaction samples used in this work, and J. Holt for insightful discussions. The IMSRG code used in this work makes use of the Armadillo C++ library [53, 54]. Computational resources were provided by subMIT at MIT Physics. This work was supported by the Office of Nuclear Physics, U.S. Department of Energy, under grants DESC0021176 and DE-SC0021179. We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [PDF-587464-2024].